OpenStudy (anonymous):

What is the rate of change from x = π to x = 3 pi over 2?

OpenStudy (anonymous):

what is the y values?

OpenStudy (anonymous):

they didnt give them

OpenStudy (anonymous):

\[x = \pi\] then a change happens,which results in: \[x = \frac{ 3 \times \pi }{2 }\] The rate of change = First guess: \[\frac{ 3 \times \pi }{2 / \pi}\] \[= \frac{ 3 \times \pi }{ 2} \times \frac{ 1 }{ \pi }\ = \frac{ 3 }{ 2 }\] If in percentage then: \[\frac{ 3 }{ 2 } \times 100 = 150 % \] there's a increase of about 150% ! or if using delta summation: \[\pi - \frac{ 3 \times \pi }{2 } = 2\pi - 3\pi /2 = -\pi/2\] =\[= \frac{ -\pi }{2 } \div \pi / 1\] =\[-1 / 2\] @silver123:is this is so?..they haven't provided the value for y ..all these are my guess

OpenStudy (anonymous):

cool, thanks

OpenStudy (anonymous):

is the guesses are correct or not? is they comply with the rules?

OpenStudy (anonymous):

oh the choices were : a. - pi/6 b. - 6/pi c. pi/6 d. 6/pi

OpenStudy (anonymous):

then how it's cool ...all is result is wrong..any guess by you or correction

OpenStudy (anonymous):

i really have no idea...

OpenStudy (anonymous):

what's the topic of yar' chapter

OpenStudy (anonymous):

Trigonometric Functions with Periodic Phenomena